OpenStudy (anonymous):

Find the apparent weight of an object fully submerged in water if its volume is 2m^3 and its density is 4000kg/m^3.

OpenStudy (anonymous):

I know that apparent weight is the weight that the scale has to support. Also I know there are three forces acting upon the object: buoyancy and tension frm the scale going upwards while the weight is going downwards.

OpenStudy (anonymous):

i am just really confused about what equations to use---

OpenStudy (anonymous):

(Gravitionnal) weight = P = mg = Vdg = 2*4000*10=80 000 N

OpenStudy (anonymous):

Buoyant force = Vd(water)g=2*1000*10=20 000 N

OpenStudy (anonymous):

So the apparent weight is P-(Buoyant Force) = 80 000 - 20 000 = 60 000 N

OpenStudy (anonymous):

so for buoyant force it's always referring to the use of water's density and not the object's density?

OpenStudy (anonymous):

Yes indeed. Par le principe d'Archimède, ''Tout corps plongé dans un fluide subit de la part de ce fluide une poussée verticale, orientée vers le haut, dont la valeur est égale au poids du volume de fluide déplacé''.

OpenStudy (anonymous):

Which is french for:

OpenStudy (anonymous):

A body immersed in a fluid receives from the fluid a vertical thrust, facing upward, whose magnitude equals the weight of the displaced fluid.

OpenStudy (anonymous):

In this case, the fluid is water, so the the buoyancy thrust involves the density of water :)

OpenStudy (anonymous):

oh--- that makes sense. Thank you so much! ^_^

OpenStudy (anonymous):

No problem :)